40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 2 18:35:47 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
    Reply  Tue Feb 2 18:54:02 2016, Koji, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
       Reply  Wed Feb 3 23:23:47 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table KnifeEdgeScans_x.pdfBeamscan_x.pdf
          Reply  Thu Feb 4 09:16:46 2016, Koji, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
             Reply  Fri Feb 5 00:23:01 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table Beamscan_setup.pdfBeamscan_x.pdfBeamscan_y.pdfZscan.pdf
                Reply  Fri Feb 5 16:50:24 2016, gautam, Update, Green Locking, First pass at mode-matching Modematch_AUXx.pdfNewSetUp.pngModematch_AUXx_2.pdfXendModeMatch.m.zip
                   Reply  Mon Feb 8 15:36:37 2016, gautam, Update, Green Locking, Alternative mode-matching scheme Modematch_alternative.pdf
                      Reply  Tue Feb 9 04:37:10 2016, ericq, Update, Green Locking, Laser swap initiated 
                         Reply  Tue Feb 9 11:49:47 2016, gautam, Update, Green Locking, Laser swap initiated Innolight_beamProfile.pdf
                            Reply  Tue Feb 9 19:15:36 2016, gautam, Update, Green Locking, Laser swap - updates Modematch_X.pdfXendModeMatch.m.zip
                               Reply  Wed Feb 10 17:57:15 2016, gautam, Update, Green Locking, Laser swap - updates 
                                  Reply  Fri Feb 12 17:05:40 2016, gautam, Update, Green Locking, Laser swap - green light recovered but no flashes in the arm 
                                     Reply  Fri Feb 12 19:07:52 2016, Koji, Update, Green Locking, Laser swap - green light recovered but no flashes in the arm 
                                        Reply  Fri Feb 19 00:42:19 2016, gautam, Update, Green Locking, Laser swap - beam ellipticity from laser? IMG_6287.JPGIMG_6288.JPGIMG_6286.JPG
                                           Reply  Mon Feb 22 13:56:52 2016, gautam, Update, Green Locking, Laser swap -reflected beam from ETM aligned 
                                              Reply  Tue Feb 23 23:01:16 2016, gautam, Update, Green Locking, Laser swap - Green PDH locking AM_scan.pdfAM_scan_zoomed.pdf
                                                 Reply  Wed Feb 24 19:29:13 2016, gautam, Update, Green Locking, Laser swap - Green PDH locked 
                                                    Reply  Fri Feb 26 01:52:44 2016, gautam, Update, Green Locking, Laser swap - Green PDH OLTF X_PDH_OLTF_20160225.pdf
                                                       Reply  Mon Feb 29 17:17:26 2016, gautam, Update, Green Locking, Laser swap - still no green beatnote 
                                                          Reply  Wed Mar 2 17:42:19 2016, gautam, Update, Green Locking, Laser swap - some progress 
                                                             Reply  Thu Mar 3 01:25:50 2016, gautam, Update, Green Locking, Laser swap - 2 IR + 1 green beatnotes found IR_beat_20160303_2.pdf
                                                                Reply  Fri Mar 4 01:11:41 2016, gautam, Update, Green Locking, Laser swap - both green beatnotes found IR_beat_20160303_green.pdf
                                                                   Reply  Sat Mar 5 23:31:01 2016, gautam, Update, Green Locking, Laser swap - some updates AUXxTelescope.png.pngPM_setup.pdfPLLolg.pdfAMPM20160303.pdfIRbeat_20160304.pdf
                                                                      Reply  Mon Mar 7 23:51:36 2016, gautam, Update, Green Locking, Laser swap - some improvement PomonaTF.pdfXPDH.pdfYPDH.pdfgreenbeat_20160307.pdf
                                                                         Reply  Tue Mar 8 18:22:20 2016, rana, Update, Green Locking, Laser swap - some improvement 
Message ID: 12016     Entry time: Wed Mar 2 17:42:19 2016     In reply to: 12013     Reply to this: 12017
Author: gautam 
Type: Update 
Category: Green Locking 
Subject: Laser swap - some progress 

[Koji, Johannes, gautam]

With Koji's and Johannes' help, I managed to resolve the coupling the pick-off IR beam into the fiber at the X end. I will put up a more detailed elog about how this was done - but in summary, we have about 31% coupling efficiency into the fiber, which isn't stellar, but I felt this was adequate to find a beatnote. Koji also pointed out that the collimation telescope attached to the fiber at the X-end is poorly mounted - this is something to fix when we swap endtables, but this was not addressed right now because if we were to adjust this, we would also have to adjust the mode matching into the fiber.

I then attempted to tune the temperature to find the IR beatnote. While doing so, I noticed some strange features of the controller - there are essentially two display modes relevant to laser crystal temperature, one which allows us to change the setpoint and one which is an actual readback of the temperature (this one can't be adjusted). While tuning the temperature, I noticed that the latter display ("LT") did not change in value. On a hunch, I disconnected the "SLOW" control BNC on the front panel, and voila, I was able to tune the setpoint and observe the measured temperature shift accordingly. I was thus able to find a reasonably strong IR beatnote (-9dBm) at T ~ 44.6 deg C (the beat PD was set to 0dB attenuation, i.e. high gain mode). However, the moment I reconnected the SLOW control BNC, the beatnote vanished (it gradually shifted out of range of the HP network analyzer), and the same thing happens if I terminate the SLOW control BNC connector! I don't understand this behaviour, as the manual says that the range of voltages accepted to this input is +/-10V, so I would assume 0V means do nothing, but clearly this isn't the case, as the beatnote is being shifted in frequency by > 1GHz, and the tuning coefficient is listed as 5GHz/V in the manual. This situation needs further investigation.

Since I had a reasonable IR beatnote setup, I returned the HP analyzer to the control room and tried to see if a green beatnote was present as well - I first ran ASS, then maximized the green transmission using the PZT mirrors, but no beatnote is evident. The contrast isn't great, the ratio of AUX power to PSL power on the green beat PD is something like 5:1, so this probably requires some tuning as well. I will update this elog after today evening's activities...

ELOG V3.1.3-